Three runners each have a single digit number on their shirts; 1, 3 and 6. Arrange them in a row to form a three digit number that is divisible by 7.
Found 2 solutions by richard1234, Edwin McCravy:
Answer by richard1234(7193)

None of the six three-digit numbers that can be formed are divisible by 7.

Answer by Edwin McCravy(20054)

You can't, if they stand upright, because these six are the 
only possaibilities:

136 = 2×2×17, not divisible by 7
163 = a prime number, not divisible by 7
316 = 2×2×79, not divisible by 7
361 = 19×19, not divisible by 7
613 = a prime number, not divisible by 7
631 = a prime number, not divisible by 7.

However if the runner with the 6 stands on his or her head,
you'll have  931 which is divisible by 7 since 7×133=931.

Edwin